Suppose that \\( \\alpha=2i-3j+k\\) and \\(\\beta=7i-5j+k\\) , find a unit vector perpendicular to \\(\\alpha\\) and \\(\\beta) respectively.




1
Expert's answer
2021-05-23T19:36:02-0400

"\alpha = 2i-3j+k \\
\beta = 7i-5j+k"

Let "d = \alpha \times \beta"

Consider the computation

"d= \begin{vmatrix} i&j&k\\ 2&-3&1\\7&-5&1
\end{vmatrix} = (-3+5)i -(2-7)j +(-10+21)k = 2i+5j+11k"

Let c be the unit vector of d."c= \dfrac{2i+5j+11k}{\sqrt{(2)^2+(5)^2+(11)^2}} = \dfrac{2i+5j+11k}{\sqrt{150}} = \dfrac{2i+5j+11k}{5\sqrt{6}}"

We claim that c is the desired unit vector perpendicular to "\alpha" and "\beta" . To see this consider

"c \cdot \alpha = \frac{1}{5\sqrt6}(2i+5j+11k) \cdot (2i-3j+k) = \frac{1}{5\sqrt6} (4-15+11) = 0"

Also, consider

"c \cdot \beta = \frac{1}{5\sqrt6}(2i+5j+11k) \cdot (7i-5j+k) = \frac{1}{5\sqrt6} (14-25+11) = 0"